Vaughn Manufacturing sells radios for $50 per unit. The fixed costs are $605000 and the variable costs are 60% of the selling price. As a result of new automated equipment, it is anticipated that fixed costs will increase by $125000 and variable costs will be 50% of the selling price. The new break-even point in units is:

a. 29200

b. 28150

c. 30250

d. 24200
Answer #1

Variable cost = 50*60% = 30

New variable cost = 50*50% = 25

Break even point = Fixed cost/Contribution margin per unit = (605000+125000)/(50-25) = 29200 Units

So answer is a) 29200